# Large Language Models for cross-language code clone detection

### v1

@@ 4. Evaluation
    binary classifier training: we consider some basic learners that we apply to the concatenation666cf. Section 3.5 - following prior work (Keller et al., 2022), we use the Absolute difference of embeddings as a concatenation method of embeddings in code fragment pairs to train a binary classification model. As learners, we consider Support Vector Machines (SVM)  (Cortes and Vapnik, 1995) and k-Nearest Neighbors (k-NN) (Cover and Hart, 1967). For SVM, we use three variants with ball_tree, kd_tree, and brute algorithms, while for k-NN, we use four variants using the linear, polynomial, and rbf kernels. The experiments are performed using 10-Fold cross-validation on each dataset.
Results. Figure 3 describes the performance (F1 score) that can be achieved when using cosine similarity thresholds for deciding on code clones. An optimal threshold selection can achieve up to 0.92, 0.97, and 0.95 scores regarding recall, precision, and F1 for samples associated with the XLCoST dataset. For the CodeNet dataset, the scores are 0.83, 0.68, and 0.75 for recall, precision, and F1, respectively.
Figure 4 summarizes the evaluation results of the trained classifiers. SVM achieves the highest scores at 0.99 for precision, recall and F1. We further observe that k-NN achieves smilar performance to SVM on the CodeNet dataset. This suggests homogeneity in the code samples enabling the learner to have a clear decision boundary. In contrast, k-NN lags well behind SVM on the XLCoST dataset, suggesting higher diversity of samples (albeit potentially simple programs).
[RQ-4] ☛ The Text-embedding-Ada-002 embedding model yields comprehensive representations of cross-lingual code fragments that enable rapidly identifying code clones based on straightforward similarity computation or learned classification. Compared to the performance achieved by LLMs instructed by prompts of varying complexity, we find that simple models achieve better results. This suggests that a major challenge in cross-lingual code clone detection is identifying a single representation space for all programming languages rather than attempting to have a universal reasoning engine.
@@ 5. Discussion
This section explores the implications of our findings for code clone detection. Our investigation into the ability of LLMs to interpret cross-language code clones yielded valuable insights.
LLM interpretation and performance. Our experiments showed that LLMs can perform substantially better when we give them specific instructions about what to look for in the code (e.g., with the improved simple prompt) compared to when the prompt is generic and zero-shot based. The improved prompt, which instruct the LLM to reason about the overall structure and logic of the code leads LLMs to identify a large number of cross-lingual code clones (cf. Table 4). This finding suggests that LLMs can avoid dramatic reasoning failures when the prompt is sufficiently informative on the task: LLMs may need some guidance to see the bigger picture.
@@
Limitations and Threats to validity. ① This study evaluated only four LLMs and one embedding model. Expanding the research to include a broader range of models with different architectures and training data could provide a more comprehensive understanding of LLM effectiveness in this task. ② The dataset used for training and evaluation consisted of 6,000 samples per dataset. Investigating the impact of larger and more diverse datasets on LLM performance and embedding model training could be highly beneficial. ③ Further exploration of prompt design strategies is needed. This could involve investigating variations of Prompt 2, exploring entirely different prompt structures, and leveraging insights from this study to refine prompts for more optimal LLM performance. ④ Due to limitations in current LLM capabilities, most prompt approaches did not always provide definitive ”yes/no” responses for clone detection. This necessitated manual verification and correction of results, which constitute an internal threat to validity. Future research can focus on developing prompt design or LLM training strategies that yield more reliable and automated outputs. ⑤ The Text-embedding-Ada-002 model is likely based on an earlier architecture of the GPT family. Newer text embedding models may help achieve even better performance.
@@ 7. Conclusion
This study explored the potential of LLMs for cross-language code clone detection. We evaluated four LLMs and investigated how prompt engineering enhances their performance. Additionally, we examined the effectiveness of code representations in the identification of code clones. Overall, our findings suggest the potential of using powerful LLMs for cross-language code clone detection. Nevertheless, While LLMs show promise, our results suggest that simpler binary classifiers leveraging robust language-agnostic representations can achieve higher, or at least comparable, performance for cross-language code clone detection.

### v2

@@ 4. Evaluation
    Baselines model training for downstream classification task: We conducted a series of experiments using two baseline models (GraphCodeBERT and UniXcoder). Given the input limitations of the baseline models, we made some adjustments to overcome the input length constraints of the baseline models when processing the dataset. To ensure robust evaluation, the selected dataset was split into an 80% training set and a 20% validation set. This experimental setup allowed us to fine-tune the baseline models and assess their ability to identify cross-lingual code clones under comparable conditions to the LLMs.
  3. (3)
    binary classifier training: we consider some basic learners that we apply to the concatenation777cf. Section 3.5 - following prior work (Keller et al., 2022), we use the Absolute difference of embeddings as a concatenation method of embeddings in code fragment pairs to train a binary classification model. As learners, we consider Support Vector Machines (SVM)  (Cortes and Vapnik, 1995) and k-Nearest Neighbors (k-NN) (Cover and Hart, 1967). For SVM, we use three variants with ball_tree, kd_tree, and brute algorithms, while for k-NN, we use four variants using the linear, polynomial, and rbf kernels. The experiments are performed using 10-Fold cross-validation on each dataset.
Results. Figure 3 describes the performance (F1 score) that can be achieved when using cosine similarity thresholds for deciding on code clones. The optimal threshold set at 5 permit achieve up to 0.99, 0.97, and 0.98 scores regarding the overall recall, precision, and F1 for samples associated with the XLCoST dataset. For the CodeNet dataset, the scores are 0.79, 0.84, and 0.78 for the overall recall, precision, and F1, respectively.
Table 6 show the performance of the baseline models. They demonstrated superior performance compared to the LLMs, achieving an improvement of 15 percentage points in terms of the overall F1-score.
Table 6. Comparison of LLM, Baselines, and Binary Classifier Performance
CodeNet Models Name Clone type Recall Precision F1-score Baselines GraphCodeBERT Non-clone 0.96 0.97 0.97 Clone 0.98 0.97 0.97 Average 0.97 0.97 0.97 UnixCoder Non-clone 0.94 0.88 0.91 Clone 0.89 0.95 0.92 Average 0.92 0.92 0.92 LLM GPT-3.5-Turbo Non-clone 1.00 0.75 0.85 Clone 0.66 1.00 0.79 Average 0.83 0.88 0.82 Binary classifier Text-embedding-3-large + SVM(kernel=‘poly’) Non-clone 0.99 0.98 0.99 Clone 0.99 0.99 0.99 Average 0.99 0.98 0.99
Figure 4 summarizes the evaluation results of the trained classifiers. SVM achieves the highest scores at 1.00 for precision, recall, and F1. We further observe that k-NN achieves similar performance to SVM on the CodeNet dataset. This suggests homogeneity in the code samples enabling the learner to have a clear decision boundary. In contrast, k-NN lags well behind SVM on the XLCoST dataset, suggesting higher diversity of samples (albeit potentially simple programs).
[RQ-4.1] ☛ The Text-embedding-3-large embedding model yields comprehensive representations of cross-lingual code fragments that enable rapidly identifying code clones based on straightforward similarity computation or learned classification.
To provide a comprehensive understanding of our results, we present in table 6 an in-depth comparison between traditional baselines, LLMs, and binary classifiers.
[RQ-4.2] ☛ Compared to the performance achieved by baselines (GraphCodeBERT and UnixCoder), and LLMs instructed by prompts of varying complexity, we find that simple models (binary classifiers) achieve better results. This suggests that a major challenge in cross-lingual code clone detection is identifying a single representation space for all programming languages rather than attempting to have a universal reasoning engine.
@@ 5. Discussion
This section explores the implications of our findings for cross-language code clone detection. Our investigation into the ability of LLMs to interpret cross-language code clones yielded valuable insights.
LLM interpretation and performance. Our experiments showed that LLMs can perform substantially better when we give them specific instructions about what to look for in the code (e.g., with the improved simple prompt) compared to when the prompt is generic and zero-shot based. The improved prompt, which instructs the LLM to reason about the overall structure and logic of the code leads LLMs to identify a large number of cross-lingual code clones (cf. Table 4). This finding suggests that LLMs can avoid dramatic reasoning failures when the prompt is sufficiently informative on the task: LLMs may need some guidance to see the bigger picture.
@@
Potential Data leakage. Our experimental results using XLCoST demonstrate that the GPT-3.5-turbo exhibits a notable degree of familiarity with the samples. This may suggest potential leakage within this LLM training set.
Limitations and Threats to validity. ① This study evaluated only five LLMs and three embedding models. Expanding the research to include a broader range of models with different architectures and training data could provide a more comprehensive understanding of LLM effectiveness in this task. ② The data used in our evaluations are derived from widely used datasets in the literature. However, they may lack diversity, and this can lead to different observations on more diverse data (in terms of programming languages, complexity, …). ③ Further exploration of prompt design strategies is needed. This could involve investigating variations of Prompt 2, exploring entirely different prompt structures, and leveraging insights from this study to refine prompts for more optimal LLM performance. ④ Due to limitations in current LLM capabilities, most prompt approaches did not always provide definitive ”yes/no” responses for clone detection. This necessitated manual verification and correction of results, which constitute an internal threat to validity. Future research can focus on developing prompt design or LLM training strategies that yield more reliable and automated outputs. ⑤ The Text-embedding-3-large model is likely based on an earlier architecture of the GPT family. Newer text embedding models may help achieve even better performance. ⑦ This study assumes a balanced dataset with an equal number of positive and negative pairs, which does not align with real-world conditions. In practical software projects, the vast majority of method pairs are not clones, resulting in a highly imbalanced scenario. While the balanced dataset used in this study provides insights into the techniques’ effectiveness under controlled conditions, it may not fully represent their performance in real-world contexts. Future work should evaluate the studied techniques on datasets with more realistic imbalances to better understand their behavior and effectiveness in such scenarios. ⑧ The construction of negative pairs in the evaluation dataset introduces a potential threat to validity. While the authors employ an approach by distinguishing at the problem level to identify negative pairs, this design choice may inadvertently exclude certain types of negative pairs, particularly ”close misses” pairs of code snippets that are similar but originate from distinct problems. As a result, the dataset may not fully capture the complexity of real-world cross-lingual clone detection scenarios, where such challenging distinctions are common. Future work could address this by exploring alternative sampling strategies to include a broader variety of ”close misses” negative pairs, and by examining how different definitions of ”negative pairs” influence the evaluation outcomes. This would help create a dataset that more comprehensively reflects the diversity of challenges in real-world settings.
@@ 7. Conclusion
This study explored the potential of LLMs for cross-language code clone detection. We evaluated five LLMs and investigated how prompt engineering enhances their performance. Additionally, we examined the effectiveness of code representations in the identification of code clones. Overall, our findings suggest the potential of using powerful LLMs for cross-language code clone detection. Nevertheless, While LLMs show promise, our results suggest that simpler binary classifiers leveraging robust language-agnostic representations can achieve higher, or at least comparable, performance for cross-language code clone detection. The insights gained from this study provide a foundation for future research aimed at developing more robust LLM-based techniques and improving code representation methods to advance software engineering practices. Furthermore, the methodologies and approaches presented here offer valuable reference points for continued exploration in this emerging domain. Future work could focus on addressing dataset imbalances by simulating ”close misses” or exploring alternative sampling strategies to construct more representative negative pairs, thereby enhancing the reliability and applicability of cross-language code clone detection techniques.
Data Availability
To promote transparency and facilitate reproducibility, we make our artifacts available to the community at:
https://anonymous.4open.science/r/clccd/
The repository includes the constructed benchmark of cross-lingual code clones, the prompt details, the experiment scripts, and the results.